# A Graph Neural Network Approach to Dispersed Systems

### v1

@@ III.1 Governing equations
Here the symbols stand for the velocity of the particles, $\boldsymbol{v}=\begin{bmatrix}\boldsymbol{v}_{1},\boldsymbol{v}_{2},\dots\end{bmatrix}^{{}^{\intercal}}\in\mathbb{R}^{N\times 2}$; position of the particles, $\boldsymbol{x}=\begin{bmatrix}\boldsymbol{x}_{1},\boldsymbol{x}_{2},\dots\end{bmatrix}^{{}^{\intercal}}\in\mathbb{R}^{N\times 2}$; vectors of, respectively, external, $\boldsymbol{F}^{\text{external}}(\boldsymbol{x})\in\mathbb{R}^{N\times 2}$, and interaction, $\boldsymbol{F}^{\text{interaction}}(\boldsymbol{x})\in\mathbb{R}^{N\times 2}$, forces; thickness, $s$, and density, $\rho$, of the particles; and time, $t$. For this ‘toy’ problem, we assume that the drag force on each particle is proportional to the particle’s velocity and radius with a constant of proportionality $k$.
@@
where $\rho_{f}$ is the density of the medium, and $g$ is the gravitational acceleration. We further consider a modified, truncated, Lennard-Jones (LJ) potential to model the inter-particle interactions. Specifically, for particles $i$ and $j$,
@@
III.2 Boundary conditions
We impose solid-wall boundary conditions with an energy loss ratio $\gamma\in(0,1]$; i.e., upon each collision to a wall, the kinetic energy of the colliding particle is reduced to $\gamma$ multiplied to its pre-collision value. For simplicity we suppose that this energy loss affects both velocity components equally, so that the post-collision velocity is scaled by $\sqrt{\gamma}$. Furthermore, the direction of the particle motion after collision is determined by assuming a mirror reflection from the wall. If we denote the velocity vectors before and after collision by $\boldsymbol{v}$ and $\boldsymbol{v}^{\prime}$, respectively, then such a reflection can be expressed as
@@ III.3 Numerical Integration
We use the Euler method to numerically integrate the equation of motion (2). Of course, we could use a more accurate numerical scheme such as RK4. But note that the main objective here is to generate some trajectories to train our GNN and put it into a test, rather than developing a comprehensive, accurate, physics-based model and simulator.
IV Results, Discussion, and Future Work
We used a total of $16$ full-length simulations each with approximately $500$ time-steps, which, for $C=6$, translates to $16\times(500-6)=7904$ training examples. We trained our model over $2$ million steps with a batch size of $2$ examples, i.e., $\approx 500$ epochs. For each simulation, a few hundred particles with randomly selected radii from a prespecified list were placed in the upper half of a box at random positions and with random initial two-dimensional velocities.
@@ IV Results, Discussion, and Future Work
Our main target for the model was a rollout loss, $\mathcal{L}_{\text{rollout}}$, defined as the mean squared error between the true and predicted positions (normalized by the box size $L$) across all time steps, particles, and spatial dimensions in a full-length test simulation.
@@
Fig. 4 shows a comparison between the ground-truth and predictions of our trained GNN for a test simulation (see Supplementary Video 1 for an animation). The trained GNN accurately predicts the particle positions over long trajectories and captures important spatial structures such as clusters and relative placements of particles. Additionally, it learns the intricate nature of particle-particle and particle-wall interactions/collisions. In particular, the particles never cross the box boundaries. It is important to note that we are not applying any mask or penalty to achieve this. The model’s accuracy slightly drops as the particles get crowded at the bottom of the box. We argue, however, that more training data will resolve this issue. The model simply has not seen enough trajectories in such relatively more complex configurations.
@@
We stress that the particle-size could be treated as a categorical property, where certain particle sizes (categories) were considered in training. As such, particle radius would act as a particle type. However, we were interested in treating the particle radius as a continuous quantity, and in developing a model that predicts the dynamics of a system with any particle size. In fact, we tested our trained model on a binary system with particle sizes not present in the training data. Fig. 5 and Supplementary Video 2 show the corresponding results. The GNN does an excellent job in predicting the particle trajectories despite not seeing these particles sizes during training. We should note however, that our GNN does not show equally accurate results for systems involving particles with radii far outside the radius range used in training. Similarly, one could use training data that include different fluids, particle materials, particle-fluid interfacial properties, etc. to develop a comprehensive model that captures the impact of various physical properties of the problem.
@@
    Inclusion of Rotation. An exciting future work is to account for the inevitable rotation of the particles in two and three dimensions. For such a problem, orientation and angular velocity of the particles contribute to the node and edge features, and the GNN predicts both the translational and rotational accelerations of the particles during training. Since this problem is inherently more complex, it would put our machine learning framework into a more rigorous test. The training data can be obtained through either experiments and high-speed-imaging, or a simulation software such as Ansys Fluent. However, the former, albeit more appealing to us, faces some challenges in a three-dimensional system (e.g., particles masking each other).
Figure 5: Comparison of the ground truth (left panel) and the GNN prediction for a system with particle sizes not present during training. $\mathcal{L}_{\text{rollout}}=6.5\times 10^{-4}$.
@@
    Asymmetric Particles. Our GNN can be extended to systems with non-spherical particles. One can represent particles with multiple geometrical characteristics instead of a mere radius that we used here for spheres (e.g., length and thickness for a rod-like particle). Additionally, relative orientation of the particles plays an important role in particle-particle interactions/collisions, and particle-wall collisions as well. While we acknowledge that this is a significantly more complex problem, we believe it is a feasible future project.
@@
    Colloidal Dispersions. Another avenue for future work is the application of the proposed GNN framework to experimental colloidal dispersions. Recent advances, such as those demonstrated by ActiveNet [18], highlight the potential of GNNs to uncover active and two-body forces in colloidal systems by learning directly from experimental trajectories. Extending our approach to incorporate these techniques could enable the study of forces and dynamics in a wider range of active and passive colloidal systems, including systems under external fields or exhibiting many-body interactions.

### v3

@@ III Sample Dataset Generation
Note that we consider the deterministic motion of the particles, with no stochastic processes such as thermal fluctuations involved. Here the symbols stand for the velocity of the particles, $\boldsymbol{v}=\begin{bmatrix}\boldsymbol{v}_{1},\boldsymbol{v}_{2},\dots\end{bmatrix}^{{}^{\intercal}}\in\mathbb{R}^{N\times 2}$; position of the particles, $\boldsymbol{x}=\begin{bmatrix}\boldsymbol{x}_{1},\boldsymbol{x}_{2},\dots\end{bmatrix}^{{}^{\intercal}}\in\mathbb{R}^{N\times 2}$; vectors of, respectively, external, $\boldsymbol{F}^{\text{external}}(\boldsymbol{x})\in\mathbb{R}^{N\times 2}$, and interaction, $\boldsymbol{F}^{\text{interaction}}(\boldsymbol{x})\in\mathbb{R}^{N\times 2}$, forces; thickness, $s$, and density, $\rho$, of the particles; and time, $t$. For this ‘toy’ problem, we assume that the drag force on each particle is proportional to the particle’s velocity and radius with a constant of proportionality $k$.
@@
where $\rho_{f}$ is the density of the medium, and $g$ is the gravitational acceleration. We further consider a modified, truncated, Lennard-Jones (LJ) potential to model the interactions between particles. Specifically, for particles $i$ and $j$,
@@
We could also consider other types of interactions, such as hydrodynamics, or electrostatics for systems involving charged particles or ions. However, the Lennard-Jones interactions used in this toy model provide sufficient complexity in terms of stiffness and spatial dependence to serve the main goal of rigorously testing the GNN. The primary focus here is not to develop a highly accurate physical model, but to generate suitable datasets for the GNN’s evaluation. In this context, the GNN will ultimately focus on learning from the data, irrespective of the specific physical model used to generate it. Nonetheless, in cases where the nature of interactions significantly differs, it may be necessary to adjust the network architecture or features to ensure effective learning.
To complete the description of the toy model, we impose solid-wall (i.e., no overlap upon collision) boundary conditions with an energy loss ratio $\gamma\in(0,1]$; i.e., upon each collision with a wall, the kinetic energy of the colliding particle is reduced to $\gamma$ times its pre-collision value. For simplicity we suppose that this energy loss affects both velocity components equally, so that the post-collision velocity is scaled by $\sqrt{\gamma}$. Furthermore, the direction of the particle motion after collision is determined by assuming a mirror reflection from the wall. Denoting the velocity vectors before and after collision as $\boldsymbol{v}$ and $\boldsymbol{v}^{\prime}$, respectively, this reflection can be expressed as
@@
To generate particle trajectories, we use the Euler method to numerically integrate the equation of motion (2). While more accurate numerical schemes could be employed, the primary objective here is to create datasets for training and testing our GNN, rather than developing a comprehensive, highly accurate, physics-based model or simulator.
For the simulations, we use the following parameter values: $s=2\;\mathrm{mm}$, $\rho=2300\;\mathrm{kg\cdot m^{-3}}$, $\rho_{f}=1000\;\mathrm{kg\cdot m^{-3}}$, $g=9.81\;\mathrm{m\cdot s^{-2}}$, $L=10\;\mathrm{cm}$, $k=1.25\;\mathrm{kg\cdot s^{-1}\cdot m^{-1}}$, and $\gamma=0.8$. We also use $\Delta t=10^{-4}\;\mathrm{s}$, but take a snapshot of the particle positions every $15$ time steps. For each simulation, a few hundred particles with randomly selected radii from a pre-specified list were placed in the upper half of the box at random positions and with random initial two-dimensional velocities. We store the positional data and particle radii scaled by the box size $L$. For further details, please refer to the corresponding code script in the repository.
IV Results & Discussion
We performed a total of $16$ full-length simulations, each generating approximately $500$ snapshots. With $C=6$, this yields $16\times(500-6)=7904$ training examples. The model was trained for $2$ million steps with a batch size of $2$ examples, equivalent to approximately $500$ epochs.
@@ IV Results & Discussion
Recall that the loss function (1) compares the true and predicted accelerations of the particles, given a sequence of last $C$ positions. Our ultimate goal, however, is to make accurate rollout predictions, i.e., full trajectories of all particles over time. To quantify the error in our rollout predictions, we introduce a rollout loss, $\mathcal{L}_{\text{rollout}}$, defined as the mean squared error between the true and predicted positions (normalized by the box size $L$), across all time steps, particles, and spatial dimensions in a full-length test simulation.
@@
Fig. 4 compares the ground-truth and the predictions of our trained GNN for a test simulation (see Supplementary Video 1 for an animation). The trained GNN accurately predicts the particle positions over long trajectories and captures important spatial structures, such as clusters and relative placements of particles. Furthermore, it learns the intricate nature of particle-particle and particle-wall interactions and collisions. Notably, the particles never cross the box boundaries, even though no mask or penalty is applied to enforce this behavior. The model’s accuracy slightly drops as the particles get crowded at the bottom of the box. We argue, however, that more training data will resolve this issue. The model simply has not seen enough trajectories in such relatively more complex configurations. The rollout loss for six randomly generated test cases had a mean value of $9.5\times 10^{-4}$ with a standard deviation of $2.2\times 10^{-4}$. For the results shown in Fig. 4, $\mathcal{L}_{\text{rollout}}=6.6\times 10^{-4}$. To provide insight into the physical significance of this loss, for $L=10\;\mathrm{cm}$, it translates to an average discrepancy of $0.066\;\mathrm{mm}$ between the ground-truth and predicted trajectories. For an average particle diameter of $6\;\mathrm{mm}$, this results in a mere $\approx 1\%$ error relative to the particle size. Also, visually, it is quite challenging to distinguish the predicted trajectories from the ground truth, further underscoring the accuracy of the model.
@@
We emphasize that the particle size could be treated as a categorical property, where distinct particle sizes (categories) are considered during training, and the particle radius would act as a particle type. However, we were interested in treating the particle radius as a continuous quantity, and in developing a model that predicts the dynamics of a system with any particle size. To demonstrate this point, we tested our trained model on a system with particle sizes not included in the training data. Fig. 5 and Supplementary Video 2 show the corresponding results. Remarkably, the GNN accurately predicts the particle trajectories despite not having seen these particle sizes during training. For $12$ test cases involving only particle sizes not present during training, the average rollout loss was $9.4\times 10^{-4}$ with a standard deviation of $2\times 10^{-4}$. This performance is noteworthy, as the GNN generates predictions for previously unseen particle sizes with a loss comparable to the predictions for particle sizes seen during training. We should note, however, that the model’s accuracy diminishes for systems involving particles with radii far outside the radius range used in training. Similarly, one could use training data that include different fluids, particle materials, particle-fluid interfacial properties, etc. to develop a comprehensive model that captures the impact of various physical properties on the system.
Figure 5: Comparison of the ground truth (left panel) and the GNN prediction (right panel) for a system with particle sizes not present during training. $\mathcal{L}_{\text{rollout}}=6.5\times 10^{-4}$.
V Conclusions
@@
    Inclusion of Rotation. An exciting future work is to account for the inevitable rotation of the particles in two and three dimensions. In this scenario, the orientation and angular velocity of the particles contribute to the node and edge features, and the GNN predicts both the translational and rotational accelerations of the particles during training. Since this problem is inherently more complex, it would put our machine learning framework into a more rigorous test. The training data can be obtained either through experiments and high-speed-imaging, or using simulation software such as Ansys Fluent. However, the former, albeit more appealing to us, faces some challenges in a three-dimensional system (e.g., particles masking each other).
@@
    Asymmetric Particles. Our GNN can be extended to systems with non-spherical particles. One could represent particles with multiple geometrical characteristics, rather than just a single radius as we used here for spheres (e.g., length and thickness for a rod-like particle). Additionally, the relative orientation of the particles plays an important role in the particle-particle and particle-wall interactions and collisions. While we acknowledge that this is a significantly more complex problem, we believe it is a feasible future project.
@@
    Colloidal Dispersions. Another avenue for future work is the application of the proposed GNN framework to experimental colloidal dispersions. Recent advances, such as those demonstrated by ActiveNet [36], highlight the potential of GNNs to uncover active and two-body forces in colloidal systems by learning directly from experimental trajectories. Extending our approach to incorporate these techniques could enable the study of forces and dynamics in a wider range of active and passive colloidal systems, including those under external fields or exhibiting many-body interactions.
Acknowledgments
We thank Dr. A. Molaei Dehkordi for insightful discussions on the time complexity of computer simulations of fluidized beds.
Statements and Declarations
Competing Interests. There are no conflicts to declare.
Data Availability. All of the codes along with the training and test data used to write this paper can be found on GitHub at github.com/rfjd/GNS-DispersedSystems.